OpenStudy (anonymous):

if h(x)=g(x)(f(x)+1), where f(2)=-3, g(2)=4, f'(2)=-2, and g'(2)=7, find h'(2). this is what i have: h'(2)=g'(2)(f(2)+1)+(f(2)+1)'g(2) h'(2)=7(-3+1)+(-2)4 h'(2)=-22 is this right?

OpenStudy (anonymous):

i say yes your right

OpenStudy (anonymous):

how sure are you?

TheSmartOne (thesmartone):

Yes there is @NH4C But you can just post in the main Math area :)

OpenStudy (jhannybean):

For specific calculus - related problems, post in the calculus section. Then you will get help from people who know how to solve calculus-based problems and faster solutions than people who only take stabs at the problem.

OpenStudy (unklerhaukus):

h(x) = g(x)(f(x)+1) h(2) = g(2)(f(2)+1) h'(2) = g'(2)(f(2)+1)+(f(2)+1)'g(2) h'(2) = 7(-3+1)+(-2)'4 h'(2) = 7(-2)+0(4) derivative of a constant is zero h'(2) = -14